### v2

# Expansion-contraction duality breaking in a Planck-scale sensitive cosmological quantum simulator

Abstract
We propose the experimental simulation of cosmological perturbations governed by a Planck-scale induced Lorentz violating dispersion, aimed at distinguishing between early-universe models with similar power spectra. Employing a novel variant of the scaling approach for the evolution of a Bose-Einstein condensate with both contact and dipolar interactions, we show that scale invariance, and in turn, the duality of the power spectrum is broken at large momenta for an inflating gas, and at small momenta for a contracting gas. We thereby furnish a Planck-scale sensitive approach to analogue quantum cosmology that can readily be implemented in the quantum gas laboratory.
Inflation [1, 2, 3, 4, 5, 6] provides a causal mechanism for the generation of primordial density perturbations with a nearly scale-invariant power spectrum as observed in the Cosmic Microwave Background (CMB) [7, 8]. However, a key limitation arises from the possibility that trans-Planckian modes generated close to the initial singularity could also have redshifted to observable scales. While a self-consistent treatment of such modes would require a UV-complete framework, efforts to test the robustness of Hawking radiation against modified dispersions (arising e.g. in black-hole analogs [9]) inspired ad hoc models of trans-Planckian physics [10, 11, 12, 13]. In the cosmological context, such models revealed that scale-invariance was generally not robust to short-distance modifications [14, 15, 16, 17, 18, 19, 20, 21, 22, 23] – tightly constraining the assumptions on trans-Planckian physics in inflationary scenarios [24, 25, 26, 27, 28].
The idea of a bouncing cosmology [29] can circumvent the initial singularity and prevent trans-Planckian modes from reaching observable scales: An initial contraction phase generates primordial density perturbations (as schematically illustrated in 1) with a scale-invariant power spectrum similar to inflation via a duality invariance of perturbation spectra corresponding to certain expanding and contracting backgrounds [30, 31, 32]. Since both inflation [33, 34] and bouncing [35, 36, 37] models encounter unresolved conceptual (as well as fine-tuning) issues, this duality in fact weakens the power spectrum as a unique indicator of early-universe possibilities [38, 39].
Given the inherent challenges in recreating the initial conditions of the Universe, cosmology has largely relied on observations. To enable addressing cosmological issues in reproducible experiments, the era of analogue quantum cosmology using ultracold gases was ushered in, first theoretically  [40, 41, 42, 43, 44, 45], and then culminating in pioneering experiments [46, 47, 48, 49, 50]. For reviews of what has more generally, covering a broad range of physical systems, been dubbed analogue gravity, see [51, 52, 53]. Analogue gravity has led to major milestones, most notably the first observation of the quantum Hawking radiation effect in a Bose-Einstein condensate (BEC) [54, 55, 56]. These systems also allow, in principle, the observation of the more elusive quantum Unruh effect in its various guises [57, 58]. While contact interactions between the gas constituents have remained the primary focus, recent studies taking into account dipole-dipole interactions [59] have significantly enriched BEC simulations, in particular for exploring the impact of trans-Planckian dispersion cf., e.g., Refs. [60, 61, 62].
Figure 1: Timeline of comoving horizon $|aH|^{-1}=|\frac{\eta}{v}|$ (corresponding to cosmological scale factor $a\propto|\eta|^{v}$), and comoving mode propagation prior to reaching current observable CMB scales. A scale-invariant power spectrum can be generated in the early epoch ($\eta<0$), either via inflation ($v=-1$), or via a contraction phase ($v=3$) leading up to bounce (set at $\eta=0$).
In what follows, we propose the experimental simulation of primordial density perturbations in a quasi-2D dipolar BEC, wherein a strongly confining transverse trap introduces an effective Planck-scale that dynamically alters the standard Lorentz-invariant dispersion. We show that a novel anisotropic variant of the usual scaling approach [63, 64] can be used to engineer a dispersion of the form $\omega_{k}\propto a\lambda_{pl}^{-1}F[k\lambda_{pl}/a]$, previously employed to address trans-Planckian issues in cosmology [14, 15, 16, 17, 18, 19, 20, 21, 22, 23], and which arises in UV-complete quantum gravity candidates (such as Horava-Lifshitz [65, 66, 67]) as well as string-theory motivated minimal-length models [68, 69]. It is explicitly demonstrated that the duality invariance of the power spectrum corresponding to expanding–contracting cosmologies [30] is broken by trans-Planckian physics, as can be experimentally verified in our quantum simulator. Specifically, emergent nonadiabatic corrections tilt the spectrum at large momenta for inflation and at small momenta for contraction. Aided by these observations, we discuss how cold-atom experiments can isolate potential large-scale signatures of Lorentz violation in the power spectrum, and distinguish between competing early-universe models.
We further highlight a microscopically controlled variant of the dispersion first put forward by Unruh [11]. This “quasi-flat band” (nearly zero group velocity over a large range of wavevectors) dispersion is shown to result in the freezing (akin to inflation) of an otherwise growing contraction power spectrum, along with a slight red-tilt similar to current CMB observations [70]. Remarkably, the Unruh dispersion corresponds to exactly equal dipolar ($g_{d}$) and contact ($g_{c}$) couplings, which coincides with the stability boundary of the bulk quantum gas in the embedding three spatial dimensions of our 2+1D setup, cf. Refs. [71, 72]. The Unruh dispersion is, thus, readily experimentally realizable in dipolar-contact BECs with only moderate Feshbach tuning of the contact interaction necessary, for example in Dysprosium and Erbium condensates with their large magnetic dipole moments of $10$ and $7\,\mu_{B}$, respectively [73, 74]. Hence Planck-scale sensitive analogue cosmology can be realized with current experiments in the quantum gas laboratory.
Setup. We consider a 3D quantum gas subject to contact as well as dipole-dipole interactions, with the dipoles aligned along the $z$ direction. We confine the gas to a harmonic oscillator ground state in the transverse direction [71] with an evolving trap frequency $\omega_{z}(t)$ that scales the oscillator length as $d_{z}(t)=b_{z}(t)d_{z,0}$, and integrate out the $z$ direction; thus leaving behind an effectively quasi-2D condensate. We then employ the scaling approach — an established procedure for analyzing BECs with generally time dependent coupling strengths placed in time-varying traps [63, 64]. However, we relax a central assumption in the otherwise general scaling approach of [75], namely that the time dependence of pairwise interaction terms can be collected as $V(\mathbf{r};t)=\mathscr{V}(t)V(\mathbf{r})$. Distinct from earlier works, we can therefore address a form of anisotropic scaling along radial and transverse directions tailored to our simulation purposes, which results in a time dependent modified dispersion in the comoving frame of the quasi-2D condensate (for details, see Supplemental Material [76]).
In the comoving frame defined by the 2D scaled coordinate $\mathbf{x}=\mathbf{r}/b(t)$, the fluid density is approximately constant ($\rho\sim\rho_{0}$), and the velocity vanishes ($\partial_{t}\mathbf{x}\sim 0$). By synchronizing the transverse scaling with respect to the desired cosmological scale factor as $b_{z}=a^{2}$ (which allows the contact and dipolar coupling strengths, i.e., $g_{c,0}$ and $g_{d,0}$ respectively, to be kept constant [76]), and linearizing the fluctuations on top of the condensate phase ($\phi_{0}+\delta\phi$) and density ($\rho_{0}+\delta\rho$), we get the corresponding equations of motion in the momentum space as follows:
$\delta\ddot{\phi}_{k}$ $+\left(2\frac{\dot{a}}{a}-\frac{\dot{W}_{k}}{W_{k}}\right)\delta\dot{\phi}_{k}+\frac{c_{0}^{2}k^{2}W_{k}}{a^{2}}\delta\phi_{k}=0,$
$\delta\ddot{\rho}_{k}$ $+\frac{c_{0}^{2}k^{2}W_{k}}{a^{2}}\delta\rho_{k}=0,$ (1)
where we have defined, cf. [71]
$W_{k}$ $=1-\frac{3R}{2}w\left[\frac{b_{z}kd_{z,0}}{b}\right]+\frac{k^{2}d_{z,0}^{2}a^{2}}{4A};$ (2)
$R$ $=\frac{g_{d,0}}{d_{z,0}g_{\rm eff,0}};\quad A=\frac{mc_{0}^{2}}{\hbar\omega_{z,0}};\quad c_{0}^{2}=\frac{g_{\rm eff,0}\rho_{0}}{m},$
[…]
$\square\delta\phi$ $=\frac{1}{\sqrt{|g|}{}}\partial_{\mu}\left(\sqrt{|g|}g^{\mu\nu}\partial_{\nu}\delta\phi\right)=0;$
$ds^{2}$ $=g_{\mu\nu}dx^{\mu}dx^{\nu}=d\tau^{2}-a^{2}(\tau)d\mathbf{x}^{2}.$ (3)
The second and third terms in $W_{k}$ correspond to dipolar interaction and free-particle contribution respectively, which break Lorentz invariance at shorter wavelengths.
By appropriately tuning the parameters $R$ (relative dipolar strength) and $A$ (dimensionless sound speed), one can simulate a variety of nonlinear trans-Planckian dispersions, as shown in 2. However, our aim here is to probe Planck-scale effects in cosmology via a dispersion of the form $k^{2}W_{k}|_{A\to\infty}=a^{2}\lambda_{pl}^{-2}F^{2}(k\lambda_{pl}/a)$ [14, 15, 16, 17, 18, 19, 20, 21, 22, 23] — wherein high-momentum modes fall back to the standard Lorentz-invariant dispersion after being redshifted to sub-Planckian scales by the expansion. In our system, this dispersion can now be realized by further synchronizing the trapping frequencies along radial and transverse directions such that $b=ab_{z}$, leading to
$W_{k}=1-\frac{3R}{2}w\left[\frac{kd_{z,0}}{a}\right]+\frac{k^{2}d_{z,0}^{2}a^{2}}{4A},$ (4)
where the Planck length $\lambda_{pl}$ for our effective 2+1D space-time is set by the trap-width $d_{z,0}$ along the compactified extra (transverse) dimension. We confine the relative strength $R$ up to a critical value $R_{c}\coloneqq\sqrt{2\pi}/3$, at which the dispersion coincides almost exactly with that of Unruh [11], $k^{2}W_{k}|_{A\to\infty}\approx a^{2}d_{z,0}^{-2}\tanh^{2/p}{\left[(kd_{z,0}/a)^{p}\right]}|_{p\to 1/2}$, asymptoting to zero group velocity at large momentum (2). Beyond the critical strength $R_{c}$, the dispersion becomes unstable at large momentum, unless stabilized by the free-particle term (resulting in a roton minimum [59]). Though this free-particle term resembles a Corley-Jacobson type modification (2[12], it amplifies Lorentz violation with the redshifting of modes, and hence does not model any early universe effect of relevance. We therefore confine the experimental protocol to time- and momentum-scales where the free-particle contribution can be safely ignored. Beyond these scales, the mapping to cosmological perturbation theory breaks down, see [76]. The limit of very large $A\gg 1$ (keeping $R\leq R_{c}$), henceforth assumed, thus captures particularly well the low-momentum signatures we aim to address.
Figure 2: Side-by-side comparison of dispersions corresponding to (a) well-known trans-Planckian models, and (b) quasi-2D dipolar BECs. One can simulate subluminal (e.g., Unruh [11]) or superluminal (e.g., C-J for Corley-Jacobson [12]) cases by tuning $R$ and $A$. Dispersions with a minimum (such as those appearing in generalizations of the C-J type [79]) can also be modeled via roton minimum tuning (dashed red line) [71, 60].
[…]
$\delta\bar{\phi}^{\prime\prime}_{k}+\omega_{k}^{2}\delta\bar{\phi}_{k}=0;\quad\omega_{k}^{2}=k^{2}W_{k}+\frac{v(2-v)}{4\eta^{2}}\left(1+\Delta_{k}\right);$
$\Delta_{k}=\frac{(4v-2)a\partial_{a}W_{k}+2va^{2}\partial_{a}^{2}W_{k}}{(2-v)W_{k}}-\frac{3va^{2}(\partial_{a}W_{k})^{2}}{(2-v)W_{k}^{2}},$ (5)
where we have assumed a power-law scale factor $a=(\eta/\eta_{i})^{v}$. Note that the term $\Delta_{k}$ is generally absent in ad hoc trans-Planckian models in cosmology which are only confined to particular choices for the time-dependent term $k^{2}W_{k}$ [14, 15, 16, 17]. However, additional nonadiabatic corrections are known to emerge from an underlying Planck-scale sensitive framework, such as in the form of a source term for mode evolution [80, 67] or as a correction term in the dispersion  [68, 69] (i.e., $\Delta_{k}$ in (Expansion-contraction duality breaking in a Planck-scale sensitive cosmological quantum simulator)) — our setup therefore goes well beyond the ad hoc models in capturing potential observable signatures.
To first understand the $\Delta_{k}\sim 0$ case, let us consider a nearly static $W_{k}$ which can be realized in the lab via isotropic scaling ($b_{z}=b$) of the dipolar condensate [60], away from the free-particle regime (assuming $A\gg 1$). The mode functions evolving from the (minimum energy [14, 81]) vacuum state defined at $\eta\to-\infty$ are given below:
$\delta\bar{\phi}_{k}=\frac{\sqrt{\pi\absolutevalue{\eta}}}{2}H_{s}^{(1)}\left(\omega_{k}^{\rm in}|\eta|\right);\quad\lim_{\eta\to-\infty}\delta\bar{\phi}_{k}\to\frac{e^{-i\omega_{k}^{\rm in}\eta}}{\sqrt{2\omega_{k}^{\rm in}}},$ (6)
where $\omega_{k}^{\rm in}=k\sqrt{W_{k}}$ corresponds to the initial $k$-mode frequency, and $s=\absolutevalue{v-1}/2$ is the Hankel function index. Each mode evolves from sub-Hubble to super-Hubble scales, with curvature effects taking over as they cross the horizon at $k\eta=-1$. The power spectrum at superhorizon scales (|$\absolutevalue{k\eta}\ll 1$) takes the form:
[…]
For the special case $s=1$, the spectrum is scale invariant. This is the 2D counterpart of the standard Lorentz-invariant result in 3D [30], which here has further been generalized to an adiabatic modification to the dispersion. A useful measure for the power spectrum is the scalar spectral index $n_{s}-1=d(\ln\mathscr{P}_{\delta\phi})/d\ln{k}$ ). Scale-invariance corresponds to $n_{s}=1$ (adiabatic case considered above), a blue spectral tilt when $n_{s}>1$ (more clumpiness at shorter length scales) and a red spectral tilt when $n_{s}<1$ (more clumpiness at larger length scales).
Duality invariance. Cosmological models related by the transformation $v\to 2-v$ exhibit duality invariance, wherein the dispersion (Expansion-contraction duality breaking in a Planck-scale sensitive cosmological quantum simulator) remains invariant and the perturbation spectra share the same scale dependence [30, 31]. For instance, the scale-invariant case $s=1$ corresponds to the following dispersion (when $\Delta_{k}\sim 0$):
$\omega_{k}^{2}=k^{2}W_{k}-\frac{3}{4\eta^{2}},$ (8)
which can be attributed to either a de Sitter expansion ($v=-1$), or a contraction scenario ($v=3$) characterized by an equation of state $w\coloneqq P/\rho=1/3$ that lies between a radiation dominated ($w=1/2$) and a matter dominated ($w=0$) background in 2+1-D [82] (in 3+1, the dual to de Sitter is a matter-dominated contraction [30]). These models are dual in the sense that they both generate a power spectrum that is scale invariant at superhorizon scales, with the only difference being that $\mathscr{P}_{\delta\phi}$ freezes for de Sitter (|$\absolutevalue{H}$ | is constant) and grows for contraction (with growing |$\absolutevalue{H}$ |).
In the context of the scaling approach, duality implies that a scale-invariant power spectrum can be simulated by a dipolar Bose-gas that is either expanding exponentially ($a\propto e^{H\tau}$) or contracting as a power law ($a\propto(-\tau)^{3/4}$). The latter offers a convenient alternative for reproducing scale invariance in the laboratory due to the following reasons — (i) the free-particle crossover that cuts off the analogue-cosmology mapping during expansion does not exist for contraction, and (ii) density-contrast measurements [49] can achieve better temporal resolution for a power-law evolution over an exponentially-fast evolution, allowing a more accurate extraction of the phase-fluctuation spectrum. As shown in (7), the duality is also robust to a generalized dispersion as long as $W_{k}$ is nearly static, with scale invariance persisting despite broken Lorentz-invariance.
Figure 3: Time evolution of inflation ($v=-1$) and contraction ($v=3$) power spectra $\mathscr{P}_{\delta\phi}$ for various values of relative dipolar strength $R$. The black vertical lines indicate the horizon crossing times ($|k\eta|=1$) corresponding to the low-momentum ($k=10^{-3},10^{-2}$) modes considered here. The scale-invariance duality is preserved for $R=0$, and broken for $R>0$.
Duality breaking. To simulate trans-Planckian dispersion in the early-universe, the modification $W_{k}$ must be dynamically driven by dipole-dipole interactions subject to an anisotropic scaling ($b=ab_{z}$) of the gas along transverse and radial directions (4). As a direct consequence, the nonadiabatic correction $\Delta_{k}$ breaks the duality between inflation and contraction due to its sensitivity to the power law (it is no longer invariant under $v\to 2-v$). This leads to a modified late time ($\eta\to 0^{-}$) dispersion as follows:
$\omega_{k}^{2}\simeq\begin{cases}k^{2}W_{k}-\frac{3}{4\eta^{2}}\left(1-\frac{Rk}{a}\right)&v=-1\\ k^{2}W_{k}-\frac{3}{4\eta^{2}}\left(1-\frac{32Ra^{2}}{(R_{c}-R)k^{2}}\right)&v=3\end{cases},$ (9)
where the duality breaking at superhorizon scales becomes apparent, with the leading order corrections to the adiabatic case (8) (and in turn, to scale-invariance) being prominent at large $k$ for inflation and small $k$ for contraction. While the effect is amplified with increasing $R$ in both cases, contraction leads to a different late-time behaviour very close to the critical “Unruh” value $R\to R_{c}$ (note that the $a\to 0$ and $R\to R_{c}$ limits of $\Delta_{k}$ as defined in the second line of (Expansion-contraction duality breaking in a Planck-scale sensitive cosmological quantum simulator) do not commute):
$\omega_{k}^{2}\simeq k^{2}W_{k}-\frac{3}{4\eta^{2}}\left(5+\frac{24a^{2}}{k^{2}}\right).$ (10)
This hints at drastic nonadiabatic corrections featuring an additional zeroth order, scale-independent term absent in the noncritical case (9). It is interesting to note that in the leading order, this dispersion now matches that of a $v=-3$ power-law expansion, which for a nearly static $W_{k}$ results in a frozen power spectrum $\mathscr{P}_{\delta\phi}\propto(k^{2}W_{k})^{-1}$ that is scale invariant ($n_{s}=1$) at large $k$ and red-tilted ($n_{s}<1$) at small $k$.
Figure 4: Scale dependence of the superhorizon power spectrum $P_{\delta\phi}$ corresponding to low-momentum ($k\ll 1$) modes for inflation (black line) and contraction (dashed lines). Lorentz violation tilts the spectrum at small $k$ exclusively for contraction; to a blue tilt for $0<R<R_{c}$ and a red tilt at $R_{c}\sim 0.835$.
In order to understand how these corrections translate to observable consequences, we rely on numerical simulations of the power spectrum (incorporating the exact form of $\Delta_{k}$, see for details of the simulations [76]) and track its evolution all the way to superhorizon scales $|k\eta|\ll 1$. While (9) indicates duality breaking due to nonadiabatic corrections at both the low-momentum corner (for contraction) and high-momentum corner (for inflation), our focus is on the former. From 3 and 4, we observe that in this regime ($k\ll 1$) the inflationary power spectrum remains scale invariant, whereas nonadiabatic corrections to the contraction power spectrum manifest as a blue-tilt ($n_{s}>1$) for noncritical values of $R$, and a red-tilt ($n_{s}<1$) for $R=R_{c}$. At the critical “Unruh” value $R_{c}$, the contraction power spectrum freezes (as seen in 3) due to the late-time dynamics asymptoting to that of a $v=-3$ power-law expansion (10) to leading order.
Conclusion. We have highlighted the crucial role of nonadiabaticity in encoding the cosmological power spectrum with potential low-energy imprints of Planck-scale physics. For adiabatic nonlinear dispersion ($\Delta_{k}\sim 0$), simulated via conventional isotropic scaling of dipolar condensates ($b=b_{z}$), the scale-invariance duality between inflation and contraction is preserved. When however, nonadiabatic corrections arising from a Planck-scale sensitive dispersion in our novel anisotropic scaling setup ($b=ab_{z}$) occur, the duality is broken. Whereas the inflationary power spectrum washes out any low-momentum signature of Planck-scale physics with the expansion, contraction clearly records these signatures via a spectral tilt in the low-momentum corner — implying that although trans-Planckian modes never reach observable scales in a bouncing model, Lorentz violation can still leave imprints at large scales during the initial contraction phase. The critical “Unruh” case $R=R_{c}$ further leads to a frozen contraction power spectrum akin to inflation, but with a slight red-tilt similar to current CMB observations [70]. The significance of this fact is however limited by the sensitivity of the tilt to the form of $\Delta_{k}$ — which may not mimic the actual nonadiabatic correction emerging from a UV-complete theory of quantum gravity. In addition, exact matching with the CMB parameters would require extreme fine-tuning of the experimental protocol. We nevertheless stress the utility of cold-atom quantum simulators in isolating observable, low-energy effects via a Planck-scale sensitive treatment of cosmological perturbations, which facilitates distinguishing between competing early-universe models and aid our interpretation of current day observations. Finally, our findings can be validated in currently realized magnetically dipolar BECs, with only moderate Feshbach tuning of the contact interaction being applied, due to the fact that we stay outside the roton regime throughout.
This work has been supported by the National Research Foundation of Korea under Grant No. 2020R1A2C2008103.